OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL Write the rule for the following geometric sequence: 2, 10, 50, 250, …

OpenStudy (anonymous):

you got buns son

OpenStudy (anonymous):

shhh *whispers: just multiply it by five shh* XD

OpenStudy (mathstudent55):

A geometric sequence has a common ratio. A ratio is a comparison of numbers by division. Divide each number by the previous number. What do you get?

OpenStudy (anonymous):

These are my choices A. Tn= 5x2n-1 B.Tn= -5x2 n-1 C. Tn= -2x5 n-1 D. Tn= 2x5n-1

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

f u c k you guys

OpenStudy (anonymous):

@Pikachubacca